OpenStudy (anonymous):

Evaluate the function f (x) = 4 • 7x for x =-1 and x = 2. Show your work.

OpenStudy (photon336):

this means that wherever we see x we plug in x = -1

OpenStudy (photon336):

\[f(x) = 4*7x = 28x \] essentially, \[f(x) = 28x \]

OpenStudy (photon336):

now to find f(2) the value of the function when x =2 we plug in 2 for x \[f(2) = 28(2) = ? \]